OpenStudy (anonymous):

Write a system of two equations in two variables to solve the problem. How many fluid ounces of 6% sucrose solution must be mixed with 18% sucrose solution to make 900 ounces of a 10% sucrose solution?

OpenStudy (adi3):

which chapter

OpenStudy (anonymous):

not sure which chapter

OpenStudy (adi3):

okay

OpenStudy (adi3):

i'll come back later.. ned to go somewhre

OpenStudy (anonymous):

are you back?

OpenStudy (ribhu):

i can help you with this one. @mady625

OpenStudy (ribhu):

total solution would be 900 ounces

OpenStudy (ribhu):

so let 6% solution required by 'X' ounces so the requirement of 18% solution becomes 900-X

OpenStudy (ribhu):

or assume the requirement of 18 % solution be 'Y' ounces.

OpenStudy (ribhu):

since the total solution should be 900 so X+Y =900 --------------------(1)

OpenStudy (ribhu):

now final solution would contain 10% sucrose so weight of sucrose = 900*10/100 ounces = 90 ounces

OpenStudy (ribhu):

now weights of sucrose in respective solutions will be 6/100 * X & 18/100 * Y for solutions containing sucrose as 6 & 18%

OpenStudy (ribhu):

now 0.06X +0.18Y = 90 (total sucrose in final solution) ----------(2)

OpenStudy (ribhu):

these will be the two equations which will solve the above problem

OpenStudy (ribhu):

@mathrulezz @Nnesha @ganeshie8 plz see to the solution

OpenStudy (adi3):

i'm back

OpenStudy (adi3):

sorry for coming late... I was at school.... I just came now ;)